If a rectangular parking lot has a length that is 5 yards greater than the width. If the area of the parking lot is 900 square yards. Then what is the length and width.
Answer by mananth(16949)

let width be x
length = x+5
Area = L*W
x(x+5)=900
x^2+5x-900=0
x1= -5+sqrt(25+3600) /2
x1=27.6
x2=-5-sqrt(25+3600) /2
x2=-32.6
..
width = 27.6 yards
length = 32.6 yards